# Long-Context Generalization with Sparse Attention

### v2

@@ Appendix H Additional Results
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
Figure 11: Comparison of attention maps for the 2Back task. Left: In-distribution, sequence length 64. Right: Out-of-distribution, sequence length 512 (for visualization clarity, we applied max pooling with a window size 4 and stride 4). The maps are shown for the second layer for all models. We can observe that diagonal patterns are less distorted with $\alpha$-entmax. Moreover, ASEntmax mitigate dispersal of the diagonal pattern up to 3×3$\times$3 × the in-distribution sequence length and make it less distorted up to 8×8$\times$8 ×.
@@ H.2 Local Count
Figure 12: Attention maps of Entmax model on Local Count. Left: Layer 1. Right: Layer 3; We observe a local pattern: attention weights fade as relative distance increases. Input sequence: (1..1 $\times$10, 2..2 $\times$4, 3..3 $\times$10, 2..2 $\times$4, 4..4 $\times$10, 2..2 $\times$4)..$\times$4
@@ H.3 Max Retrieval
Solving this task requires an extremely concentrated attention distribution. Such distribution can be achieved by either lowering the temperature $\theta$ for softmax or increasing the entmax parameter $\alpha$. As Table 5 shows, increasing $\alpha$ yields substantial performance gains. However, if $\alpha$ becomes too large, the distribution may collapse to a one-hot vector, causing entmax to lose all gradient signal and hindering learning (e.g. $\alpha>16$). Instead, this issue can be alleviated by scaling entmax based on the sequence length. With this approach, ASEntmax with $\alpha=1.5$, learned $\beta$, and elevated $\gamma$ achieves substantially improved performance on the task.
@@ H.4 MQMTAR
We observe the same pattern across all tasks: despite theoretical extrapolation to 16×16$\times$16 ×via RoPE scaling, RoPE models poorly generalize beyond 4×4$\times$4 ×. Moreover, although models with ALiBi can extrapolate up to 64×64$\times$64 ×, ALiBi’s limited span inevitably leads to performance degradation on very long sequences. However, NAPE provides a substantial boost in all models. As in the copy task, Entmax alone underperforms Softmax, but adaptive scaling in ASEntmax makes the model superior, extending the generalization to an impressive 1024×1024$\times$1024 ×. We also conducted experiments with ASSMax to demonstrate that, despite adaptive scaling benefits and improved performance in comparison with the model without scaling, softmax dispersion still causes a significant performance drop on very long sequences.
[…]
@@ H.5 Copy
As shown by Jelassi et al. (2023), transformers generalize to the Copy task especially with appropriate positional encodings. Table 7 shows that the Softmax transformer generalizes up to 64×64$\times$64 × the ID length. Notably, SSMax outperforms all other models which might suggest that scaling is crucial for this task.
@@ H.6 Flip-Flop
We first conducted an ablation study to evaluate model performance with various RoPE scaling factors (Table H.6). Although the random baseline accuracy for Flip-Flop is $50\%$, our generative training setup with a vocabulary of 7 tokens (4 main and 3 special) can yield accuracies below $50\%$. Therefore, we treat accuracies at or below $50\%$ as poor and select a scaling factor of 16 as optimal. The RoPE scaling factor defines the expansion multiple to which the model must generalize. Throughout all experiments, however, we observe that RoPE models poorly generalize at sequence lengths 8×8$\times$8 × the in-distribution length.
@@ H.7 Reverse
From Table 11, we can see that ASEntmax with NAPE achieved impressive 8×8$\times$8 × length generalization which to our knowledge, represents the largest extrapolation reported. Moreover, RoPE models fail even at a sequence length of 96. Although NAPE improves Softmax and SSMax, it does not enable generalization beyond 1.5×1.5$\times$1.5 × the in-distribution length; however, applying adaptive scaling to Softmax (ASSMax) enables performance at 2×2$\times$2 × the in-distribution length.
@@ H.8 Sort
Table 12 demonstrates superiority of $\alpha$-entmax on Sort, with two-layer models generalizing almost perfectly to 2×2$\times$2 × under both NoPE and NAPE configurations. Furthermore, Softmax models with NAPE experience a performance decline relative to their NoPE counterparts, and adaptive scaling degrades performance for both Softmax and $\alpha$-entmax (we also report results for NoPE + SEntmax to be convinced). However, combining NAPE with adaptive scaling enhances $\alpha$-entmax. This pattern suggest that sparsity, adaptive scaling, and NAPE can act complementarily.

### v3

@@ Appendix H Detailed Results
Softmax
Refer to caption Refer to caption
Entmax $\alpha=1.5$
Refer to caption Refer to caption
ASEntmax $\alpha=1.5$
Refer to caption Refer to caption
Figure 9: Comparison of attention maps for the 2Back task. Left: In-distribution, sequence length 64. Right: Out-of-distribution, sequence length 512 (for visualization clarity, we applied max pooling with a window size 4 and stride 4). The maps are shown for the second layer for all models. We can observe that diagonal patterns are less distorted with $\alpha$-entmax. Moreover, ASEntmax mitigate dispersal of the diagonal pattern up to $3\times$ the in-distribution sequence length and make it less distorted up to $8\times$.
@@ H.2 Local Count
[…]
@@ H.3 Max Retrieval
Solving this task requires an extremely concentrated attention distribution. Such distribution can be achieved by either lowering the temperature $\theta$ for softmax or increasing the entmax parameter $\alpha$. As Table 8 shows, increasing $\alpha$ yields substantial performance gains. However, if $\alpha$ becomes too large, the distribution may collapse to a one-hot vector, causing entmax to lose all gradient signal and hindering learning (e.g. $\alpha>16$). Instead, this issue can be alleviated by scaling entmax based on the sequence length. With this approach, ASEntmax with $\alpha=1.5$, learned $\beta$, and elevated $\gamma$ achieves substantially improved performance on the task. We also note that Top-K with K=2 works roughly as well as SSMax. However, lowering K to 1 is not possible since this would essentially reduce Top-K to an argmax operation, a non-differentiable function.
@@ H.4 MQMTAR
We observe the same pattern across all tasks: despite theoretical extrapolation to $16\timesvia$ RoPE scaling, RoPE models poorly generalize beyond $4\times$. Moreover, although models with ALiBi can extrapolate up to $64\times$, ALiBi’s limited span inevitably leads to performance degradation on very long sequences. However, NAPE provides a substantial boost in all models. As in the copy task, Entmax alone underperforms Softmax, but adaptive scaling in ASEntmax makes the model superior, extending the generalization to an impressive $1024\times$. We also conducted experiments with ASSMax to demonstrate that, despite adaptive scaling benefits and improved performance in comparison with the model without scaling, softmax dispersion still causes a significant performance drop on very long sequences.
Table 9: Exact match accuracy ($\%$) on Multi-query Multi-token Associative Recall.
@@ H.5 Copy
As shown by Jelassi et al. (2023), transformers generalize to the Copy task especially with appropriate positional encodings. Table 10 shows that the Softmax transformer generalizes up to $64\times$ the ID length. Notably, SSMax outperforms all other models which might suggest that scaling is crucial for this task.
@@ H.6 Flip-Flop
We first conducted an ablation study to evaluate model performance with various RoPE scaling factors (Table H.6). Although the random baseline accuracy for Flip-Flop is $50\%$, our generative training setup with a vocabulary of 7 tokens (4 main and 3 special) can yield accuracies below $50\%$. Therefore, we treat accuracies at or below $50\%$ as poor and select a scaling factor of 16 as optimal. The RoPE scaling factor defines the expansion multiple to which the model must generalize. Throughout all experiments, however, we observe that RoPE models poorly generalize at sequence lengths $8\times$ the in-distribution length.
@@ H.7 Reverse
From Table 14, we can see that ASEntmax with NAPE achieved impressive $8\times$ length generalization which to our knowledge, represents the largest extrapolation reported. Moreover, RoPE models fail even at a sequence length of 96. Although NAPE improves Softmax and SSMax, it does not enable generalization beyond $1.5\times$ the in-distribution length; however, applying adaptive scaling to Softmax (ASSMax) enables performance at $2\times$ the in-distribution length.
@@ H.8 Sort
Table 15 demonstrates superiority of $\alpha$-entmax on Sort, with two-layer models generalizing almost perfectly to $2\times$ under both NoPE and NAPE configurations. Furthermore, Softmax models with NAPE experience a performance decline relative to their NoPE counterparts, and adaptive scaling degrades performance for both Softmax and $\alpha$-entmax (we also report results for NoPE + SEntmax to be convinced). However, combining NAPE with adaptive scaling enhances $\alpha$-entmax. This pattern suggest that sparsity, adaptive scaling, and NAPE can act complementarily.
@@
SB Attention 100.0 36.5 0 0
H.9 Language Modeling
We present a comparison between RoPE and NAPE models on standard benchmarks in Table H.9, and on length generalization in Table H.9. NAPE alone improves short-context results on LAMBADA, HellaSwag, and PIQA. Furthermore, length generalization capabilities only emerge when using either RoPE with ABF or NAPE. While SSMax, Entmax, and ASEntmax with RoPE + ABF achieve near-perfect generalization on the S-NIAH-1 task, NAPE combined with scalable models demonstrates more consistent performance across all sequence lengths in both S-NIAH-1 and S-NIAH-2. Perhaps more importantly, we find that RoPE with ABF leads to a decline in in-distribution performance. For instance, Lambada perplexity/accuracy drops, as does accuracy on HellaSwag, PIQA, ARC-C, and the in-distribution S-NIAH tasks. Moreover, both Softmax + ABF and SSMax + ABF fail dramatically on S-NIAH-2. Given that NAPE shows more consistent performance across different tasks and sequence lengths, we believe it stands out as a more robust positional encoding approach than RoPE.
Table 16: Downstream-task results on short-context datasets encompassing a comparison between RoPE and NAPE.
Method Lambada (PPL) Lambada Hellaswag PIQA Arc-C WinoGrande OpenbookQA
RoPE Softmax 59.7 30.3 32.5 64.3 26.3 50.8 28.6
SSMax 54.5 30.9 32.3 63.4 26.0 51.7 31.0
Entmax 53.0 31.1 32.9 62.9 25.8 49.9 27.8
ASEntmax 56.5 29.9 32.2 62.4 25.5 52.3 27.8
[.4pt/2pt] + ABF, Softmax 126.3 19.0 31.5 63.8 24.6 51.8 28.8
+ ABF, SSMax 96.8 21.6 31.9 63.2 24.7 51.1 30.2
+ ABF, Entmax 101.9 20.3 32.3 62.2 25.5 50.6 27.2
+ ABF, ASEntmax 99.0 20.0 29.4 62.7 24.1 52.0 29.2
[.4pt/2pt]   NAPE Softmax 52.3 30.9 33.1 65.1 25.6 49.5 28.2
SSMax 48.9 31.6 32.9 65.1 25.0 51.5 30.4
Entmax 47.9 32.1 32.8 63.6 24.6 50.9 29.0
ASEntmax 41.6 34.3 33.4 63.8 26.0 50.0 28.6
Table 17: Retrieval performance on RULER benchmark
S-NIAH-1 S-NIAH-2
ID OOD ID OOD
Model 1K 2K 4K 8K 16K 1K 2K 4K 8K
RoPE Softmax 99.8 79.0 0.0 0.0 0.0 99.6 11.4 0.0 0.0
SSMax 99.0 83.0 0.0 0.0 0.0 99.6 53.6 0.0 0.0
Entmax 100.0 79.6 0.0 0.0 0.0 99.6 53.0 0.0 0.0
ASEntmax 99.8 87.2 0.0 0.0 0.0 99.0 83.6 0.0 0.0
[.4pt/2pt] + ABF, Softmax 99.2 97.2 93.4 75.4 75.6 0.0 0.0 0.0 0.0
+ ABF, SSMax 98.2 98.0 97.6 97.4 98.0 30.8 37.4 4.4 0.2
+ ABF, Entmax 98.4 98.2 99.4 100.0 100.0 98.8 89.0 64.8 32.4
+ ABF, ASEntmax 100.0 99.6 99.6 99.6 94.0 98.6 83.6 30.8 7.2
[.4pt/2pt]   NAPE Softmax 100.0 99.4 94.2 11.4 0.8 100.0 100.0 4.8 0.0
SSMax 100.0 99.8 99.2 92.0 75.2 99.4 99.2 64.4 14.8
Entmax 99.8 99.8 89.0 21.6 1.2 99.6 99.4 64.8 7.2
ASEntmax 99.6 100.0 100.0 99.8 97.4 99.4 99.4 83.2 25.4
Appendix I AI Assistants
We used Cursor during development, and ChatGPT during paper writing for correcting grammar and polishing sentences.